OpenStudy (anonymous):

Find the exact value by using a half-angle identity. (4 points) sin 7 pie/8

myininaya (myininaya):

\[\sin(\frac{1}{2} \theta)=\pm \sqrt{\frac{1+\cos(x)}{2}}\]

myininaya (myininaya):

\[\frac{7 \pi}{8}=\frac{1}{2} \theta\] what is theta?

myininaya (myininaya):

you solve solve that and see what theta is

myininaya (myininaya):

oops x is theta

myininaya (myininaya):

\[\sin(\frac{1}{2} \theta)=\pm \sqrt{\frac{1+\cos(\theta)}{2}}\]

myininaya (myininaya):

also we know sin(7pi/8) will be positive since 7pi/8 is in the second quadrant

OpenStudy (anonymous):

mmm pie

myininaya (myininaya):

no sweets here @satellite73 only numbers that deceive you are